### v2

@@ A.1 Data Construction
We construct a dataset of West Point cadets from 1820 to 1860. We manually collect and organize information on graduates, including their Union or Confederate affiliation during the Civil War, class ranking at graduation, place of birth, year of birth, and military rank in 1865, from the Biographical Register of the Officers and Graduates of the United States Military Academy and other sources. We first present two examples of coding from main source in Figure S1.
@@
In a few cases where an individual’s year of birth was unavailable in the Biographical Register , we supplement this information by using details from Find a Grave.
@@
The information is constructed from the Biographical Register of the Officers and Graduates of the United States Military Academy. We classify as participants those who were directly engaged in combat or military operations during the Mexican-American War, as well as those who are listed as contributing indirectly to the war effort—for instance, by providing logistical support such as transporting supplies.
@@ A.3 Summary Statistics
Joining the Union (war participants) 0.115 0.320 0 1 122
Joining the Union (all) 0.066 0.249 0 1 211
Joining the war 0.578 0.495 0 1 211
Class rank 53.791 30.156 10 100 211
Age in 1860 40.652 11.273 23 65 211
Slave pop. share 45.313 9.410 33.4 57.2 211
Cohort 1841.104 11.636 1820 1860 211
1865 Career rank 5.609 2.601 2 11 207
1865 General 0.295 0.457 0 1 207
Died in civil war 0.180 0.385 0 1 211
@@
Joining the Union (war participants) 0.476 0.500 0 1 267
Joining the Union (all) 0.274 0.447 0 1 463
Joining the war 0.577 0.495 0 1 463
Class rank 54.795 28.121 10 100 463
Age in 1860 41.830 10.448 22 62 463
Slave pop. share 20.748 9.427 1.6 30.7 463
Cohort 1839.780 10.870 1820 1860 463
1865 Career rank 5.317 2.399 1 11 460
1865 General 0.241 0.428 0 1 460
Died in civil war 0.117 0.322 0 1 460
@@
Joining the Union (all) 0.515 0.500 0 1 964
Joining the war 0.560 0.497 0 1 964
Class rank 56.857 28.725 10 100 964
Age in 1860 41.527 10.931 21 63 964
Slave pop. share 0.000 0.002 0 0.01 964
Cohort 1840.354 11.364 1820 1860 964
1865 Career rank 4.711 2.135 2 10 959
1865 General 0.137 0.344 0 1 959
Died in civil war 0.091 0.288 0 1 956
Note. This table presents summary statistics for the key variables used in our analysis.
Table S3: Summary Statistics: Cohort Level
@@
Number of graduates 5.146 2.516 0 11 41
Fraction of graduates 0.127 0.056 0 0.244 41
@@
Number of graduates 11.293 3.907 3 19 41
Fraction of graduates 0.282 0.080 0.088 0.480 41
@@
Number of graduates 23.512 5.459 12 33 41
Fraction of graduates 0.591 0.084 0.462 0.794 41
@@ A.5 Peers and Personal Characteristics
joint test $p$-value 0.124 0.217 0.046 0.093 0.817 0.918 0.101 0.114
Observations 389 389 540 540 674 674 964 964
R-squared 0.016 0.016 0.013 0.013 0.002 0.002 0.006 0.006
Note. This table presents the correlations between peer composition and individual traits. After adjusting for a linear time trend (Cohort (t)), the joint test on overall correlations are statistically undistinguished from zero. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ A. Cohort level bootstrap
Share Free-State peers (sd) 0.062∗∗∗ 0.045∗∗∗ 0.040∗∗ 0.006 0.004 0.002
@@
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.220 0.268 0.000 0.015 0.028
@@ B. Cohort level cluster
Share Free-State peers (sd) 0.062∗∗∗ 0.045∗∗∗ 0.040∗∗ 0.006 0.004 0.002
@@
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.220 0.268 0.000 0.015 0.028
Note. Panel A reports the re-estimation of the baseline results using robust standard errors clustered at the cohort level. Panel B reports the corresponding re-estimates using bootstrapped standard errors clustered at the cohort level with 400 resampling iterations. (This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. ***$p<0.01$, **$p<0.05$, *$p<0.1$.)
B.2 Alternative Thresholds
@@
Thresholds: 5% Thresholds: 10%
Slave States Free States Slave States Free States
(1) (2) (3) (4) (5) (6) (7) (8)
Share Free-State peers (sd) 0.052∗∗ 0.053∗∗ -0.000 -0.004 0.059∗∗ 0.058∗∗ -0.002 -0.005
(0.024) (0.024) (0.013) (0.013) (0.023) (0.023) (0.012) (0.013)
Age in 1860 -0.010 -0.001 0.006 0.005 -0.016 -0.006 0.007 0.007
(0.016) (0.016) (0.007) (0.007) (0.018) (0.018) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.179∗∗∗ -0.039∗∗ -0.177∗∗∗ -0.070∗∗∗
(0.019) (0.019) (0.019) (0.020)
Cohort -0.012 -0.002 0.008 0.007 -0.017 -0.007 0.009 0.008
(0.015) (0.016) (0.007) (0.007) (0.017) (0.017) (0.007) (0.007)
State FEs N Y N Y N Y N Y
Dependent var. mean 0.318 0.318 0.910 0.910 0.310 0.310 0.902 0.902
Observations 352 352 577 577 339 339 590 590
R-squared 0.168 0.219 0.035 0.053 0.172 0.224 0.065 0.081
@@ B.3 Home State Economic Proxies and War Allegiances
Share Free-State peers (sd) 0.045∗∗ 0.045∗∗ 0.043∗∗ 0.042 0.004 0.003 0.003 0.003
@@
ln Farmland value per capita 0.011 0.019 0.034 0.055 0.065 0.056
@@
ln Manufactured product per capita 0.234∗∗∗ 0.097 0.009 0.021
@@
Manufacturing employment share (sd) 0.074 -0.009
@@
Dependent var. mean 0.362 0.362 0.362 0.362 0.919 0.919 0.919 0.919
Observations 389 389 389 389 540 540 540 540
R-squared 0.220 0.220 0.238 0.242 0.015 0.019 0.019 0.020
Note. This table shows the association between war allegiances and several home state economic proxies. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.4 Correlation between Academic Rank and Military Rank in 1865
Table S8: Academic rank to military rank in 1865 (The Confederacy vs. The Union)
Dependent var Rank in 1865: war participants
The Confederacy The Union
(1) (2) (3) (4) (5) (6)
Class rank 0.000 -0.001 -0.001 0.006∗∗∗ 0.006∗∗∗ 0.006∗∗
(0.003) (0.003) (0.003) (0.002) (0.002) (0.003)
Age in 1860 -0.004 -0.017 -0.034 -0.039
(0.063) (0.068) (0.043) (0.045)
Slave pop. share (sd) 0.148 0.029
(0.089) (0.063)
Cohort -0.049 -0.065 -0.087∗∗ -0.092∗∗
(0.061) (0.066) (0.042) (0.044)
State FEs N N Y N N Y
Dependent var.mean 7.516 7.516 7.516 5.848 5.848 5.848
Observations 281 281 281 637 637 637
R-squared 0.000 0.085 0.156 0.010 0.108 0.127
@@ B.5 Peer Influence by Cohort
B.6 Peer influence and shared Mexican-American War experience
Table S9: Peer composition and allegiance choice: Mexican-American War
@@ B.6 Peer influence and shared Mexican-American War experience
Slave States
Slave state cadets (>1%) Slave state cadets (1%-33%)
@@
Share Free-State peers join Mex.-Am. War (sd) $\times$ Share Free-State peers (sd) 0.246∗∗ 0.204 0.208 0.297 0.316 0.328∗∗
@@
Share Slave-State peers join Mex.-Am. War (sd) $\times$ Share Free-State peers (sd) -0.181 -0.155 -0.155 -0.227 -0.242 -0.254
(0.117) (0.124) (0.141) (0.158) (0.162) (0.156)
Share Free-State peers (sd) 0.029 0.020 0.021 0.028 0.010 0.008
(0.036) (0.033) (0.035) (0.042) (0.045) (0.045)
Share Free-State peers join Mex.-Am. War (sd) -0.152 -0.148 -0.140 -0.268 -0.247 -0.267
(0.125) (0.122) (0.136) (0.147) (0.143) (0.144)
Share Slave-State peers join Mex.-Am. War (sd) 0.129 0.120 0.120 0.237 0.221 0.244
(0.126) (0.121) (0.134) (0.145) (0.143) (0.143)
Controls N Y Y N Y Y
@@
Dependent var. mean 0.405 0.405 0.405 0.515 0.515 0.515
Observations 185 185 185 136 136 136
R-squared 0.036 0.253 0.294 0.056 0.171 0.224
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers join Mexican-American War: The proportion of Free-State peers participating in Mexican-American war with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. Share Slave-State peers joining the Mexican-American war: The proportion of Slave-State peers participating in Mexican-American War with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. Columns (1)-(3) focus on cadets from Slave States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.

### v3

@@ A.1 Data Construction
We construct a dataset of West Point cadets from 1820 to 1860. We manually collect and organize information on graduates, including their Union or Confederate affiliation during the Civil War, class ranking at graduation, place of birth, year of birth, and military rank in 1865, from the Biographical Register of the Officers and Graduates of the United States Military Academy and other sources. We first present two examples of coding from main source in Figure S1. We further supplement the dataset with additional individual- and family-level information from Ancestry.com and FamilyTree.com (and linked US Census data), including slaveholding status, parental information, spousal information, etc.
@@
In a few cases where an individual’s year of birth was unavailable in the Biographical Register , we supplement this information by using details from Find a Grave and FamilySearch.
Birth Place
The birthplace information reported in the Biographical Register of the Officers and Graduates of the United States Military Academy is occasionally incomplete or imprecise, particularly with respect to county-level identifiers. To address these limitations, we supplement the original records using genealogical sources, including Find a Grave and FamilySearch, which provide more detailed and consistent location information. These sources allow us to recover missing county-level data and to correct ambiguous or inaccurate birthplace entries, thereby improving the geographic precision of our measures.
@@
The information is constructed from the Biographical Register of the Officers and Graduates of the United States Military Academy. We record individuals who directly participated in the Mexican-American War, including those engaged in combat or other military operations.
Genealogical and Census Data Linkage
We supplemented the West Point biographical records with individual- and family-level information via manual linkage to genealogical (FamilySearch) and US Census databases (Ancestry). Using cadets’ names and birth/death dates, we first searched FamilySearch to identify the correct individual, recorded the unique page identifier, and collected birthplace, birthdate, residence around 1860, basic marital information and parental information. We then searched Ancestry for the 1860 Census and confirmed matches based on consistency in name, age, birthplace, and residence, extracting occupation, residence, and reported real and personal estate values. Finally, we consulted the 1860 Slave Schedules on Ancestry, using name and residence consistency to identify slaveholding status. Using cadets’ county of residence, we further linked each individual to county-level characteristics in 1860, including the number of churches [^15] Inter-university Consortium for Political and Social Research. United States Historical Election Returns, 1824-1968. and the voting electoral returns [^16] Haines, Michael R., and Inter-university Consortium for Political and Social Research. Historical, Demographic, Economic, and Social Data: The United States, 1790-2002..
@@ A.3 Summary Statistics
Joining the Union (War Participants) 0.107 0.310 0 1 122
Joining the Union (All) 0.062 0.242 0 1 210
Joining the War 0.581 0.495 0 1 210
Class Rank 53.810 30.068 10 100 210
Age in 1860 40.530 11.315 23 65 210
Slave Pop. Share (state level) 45.370 9.467 33.4 57.2 210
Cohort 1841.290 11.716 1820 1860 210
Join Mex.-Am. War 0.276 0.448 0 1 210
Cadet Slave Ownership 0.352 0.480 0 1 122
Slave Pop. Share (County-level) 49.415 16.553 13.191 84.999 108
Free-State Father 0.204 0.405 0 1 103
Free-State Mother 0.080 0.273 0 1 100
Free-State Wife 0.172 0.379 0 1 122
@@
Joining the Union (War Participants) 0.477 0.500 0 1 266
Joining the Union (All) 0.276 0.448 0 1 460
Joining the War 0.578 0.494 0 1 460
Class Rank 54.326 28.288 10 100 460
Age in 1860 41.841 10.461 22 62 460
Slave Pop. Share 20.739 9.379 1.6 30.7 460
Cohort 1839.761 10.890 1820 1860 460
Join Mex.-Am. War 0.317 0.466 0 1 460
Slave Ownership 0.207 0.406 0 1 266
Slave Pop. Share (County-level) 24.269 18.178 0.446 71.269 244
Free-State Father 0.132 0.340 0 1 234
Free-State Mother 0.137 0.345 0 1 233
Free-State Wife 0.297 0.458 0 1 266
@@
Joining the Union (All) 0.512 0.500 0 1 968
Joining the War 0.558 0.497 0 1 968
Class Rank 57.066 28.663 10 100 968
Age in 1860 41.535 10.899 21 63 968
Slave Pop. Share 0.000 0.002 0 0.01 968
Cohort 1840.346 11.337 1820 1860 968
Join Mex.-Am. War 0.305 0.461 0 1 968
Cadet Slave Ownership 0 0 0 0 540
Slave Pop. Share (County-level) 0.000 0.000 0 0.011 433
Free-State Father 0.868 0.338 0 1 471
Free-State Mother 0.890 0.314 0 1 462
Free-State Wife 0.467 0.499 0 1 540
Note. This table presents summary statistics for the key variables used in our analysis. Slave Pop. Share (county level) and Slave Pop. Share (state level) are expressed in percentages. Variables Free-State Father, Free-State Mother, Free-State Wife, and Cadet Slave Ownership are collected only for individuals who participated in the Civil War.
Table S3: Summary statistics: Cohort-level
@@
Number of Graduates 5.122 2.532 0 11 41
Fraction of Graduates 0.127 0.057 0 0.263 41
@@
Number of Graduates 11.220 3.863 2 19 41
Fraction of Graduates 0.280 0.080 0.059 0.480 41
@@
Number of Graduates 23.610 5.572 12 33 41
Fraction of Graduates 0.593 0.085 0.462 0.824 41
@@ A.5 Peers and Personal Characteristics
Free-State Mother 0.347 -0.124 0.347 -0.124
(0.216) (0.182) (0.216) (0.182)
Free-State Wife -0.073 0.033 -0.073 0.033
(0.141) (0.101) (0.141) (0.101)
Joint test $p$-value 0.190 0.336 0.130 0.370 0.739 0.336 0.212 0.370
Observations 388 324 540 456 670 324 968 456
R-squared 0.015 0.027 0.013 0.017 0.003 0.027 0.006 0.017
Note. This table presents the correlations between peer composition and individual traits. The joint test on overall correlations are statistically undistinguished from zero. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
A.6 Alternative Thresholds
Dependent var Join the Union: War Participants
Thresholds: 5% Thresholds: 10%
Slave States Free States Slave States Free States
(1) (2) (3) (4) (5) (6) (7) (8)
Share Free-State Peers (sd) 0.061∗∗∗ 0.065∗∗∗ -0.008 -0.011 0.064∗∗∗ 0.064∗∗∗ -0.007 -0.009
(0.022) (0.022) (0.012) (0.012) (0.024) (0.024) (0.013) (0.013)
Age in 1860 -0.008 -0.003 0.004 0.003 -0.014 -0.009 0.006 0.005
(0.015) (0.016) (0.007) (0.008) (0.016) (0.016) (0.007) (0.007)
Class Rank 0.001 0.001 -0.001∗∗ -0.001∗∗ 0.001 0.001 -0.001∗∗ -0.001∗∗
(0.001) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000) (0.000)
Slave Pop. Share (sd) -0.182∗∗∗ -0.033 -0.177∗∗∗ -0.060∗∗∗
(0.021) (0.018) (0.021) (0.020)
Cohort -0.009 -0.004 0.007 0.006 -0.015 -0.009 0.008 0.007
(0.015) (0.015) (0.007) (0.007) (0.015) (0.016) (0.007) (0.007)
State FEs N Y N Y N Y N Y
Dependent var. mean 0.317 0.317 0.911 0.911 0.308 0.308 0.905 0.905
Observations 353 353 575 575 341 341 587 587
R-squared 0.179 0.238 0.034 0.052 0.178 0.238 0.055 0.071
@@ A. Cohort level bootstrap
Share Free-State peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ -0.003 -0.005 -0.006
@@
Dependent var. mean 0.361 0.361 0.361 0.919 0.919 0.919
Observations 388 388 388 540 540 540
R-squared 0.028 0.234 0.287 0.000 0.017 0.030
@@ B. Cohort level cluster
Share Free-State peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ -0.003 -0.005 -0.006
@@
Dependent var. mean 0.361 0.361 0.361 0.919 0.919 0.919
Observations 388 388 388 540 540 540
R-squared 0.028 0.234 0.287 0.000 0.017 0.030
Note. Panel A reports the re-estimation of the baseline results using robust standard errors clustered at the cohort level. Panel B reports the corresponding re-estimates using bootstrapped standard errors clustered at the cohort level with 400 resampling iterations. (This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd), and Cohort. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. ***$p<0.01$, **$p<0.05$, *$p<0.1$.)
B.2 Different Age Groups
Table S7: Peer composition and allegiance choice: Variations in age groups
@@ B.2 Different Age Groups
Slave States Free States
(1) (2) (3) (4) (5) (6)
A. $Age<60$
Share Free-State peers (sd) 0.074∗∗∗ 0.052∗∗ 0.048∗∗ -0.002 -0.004 -0.005
(0.024) (0.021) (0.022) (0.013) (0.012) (0.012)
Controls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.359 0.359 0.359 0.920 0.920 0.920
Observations 384 384 384 535 535 535
R-squared 0.024 0.236 0.288 0.000 0.014 0.027
B. $Age<50$
Share Free-State peers (sd) 0.080∗∗∗ 0.059∗∗∗ 0.053∗∗ -0.005 -0.008 -0.009
(0.026) (0.021) (0.022) (0.013) (0.013) (0.013)
Controls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.353 0.353 0.353 0.921 0.921 0.921
Observations 334 334 334 471 471 471
R-squared 0.028 0.248 0.301 0.000 0.025 0.041
Note. Panel A restricts the sample to cadets younger than 60 at the time of Civil War participation. Panel B further restricts the sample to those younger than 50. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. ***$p<0.01$, **$p<0.05$, *$p<0.1$.)
@@ B.3 Home State Economic Proxies and War Allegiances
Share Free-State Peers (sd) 0.058∗∗∗ 0.058∗∗∗ 0.054∗∗ 0.053∗∗ -0.005 -0.006 -0.005 -0.005
@@
ln Farmland Value per Capita 0.011 0.018 0.025 0.052 0.072 0.069
@@
ln Manufactured Product per Capita 0.233∗∗∗ 0.168 0.019 0.022
@@
Manufacturing Employment Share (sd) 0.035 -0.002
@@
Dependent var. mean 0.361 0.361 0.361 0.361 0.919 0.919 0.919 0.919
Observations 388 388 388 388 540 540 540 540
R-squared 0.234 0.234 0.252 0.253 0.017 0.020 0.021 0.021
Note. This table shows the association between war allegiances and several home state economic variables. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.4 Religion and Voting
Table S9: Peer composition and allegiance choice: religion and voting
Dependent var Join the Union: War Participants
(1) (2) (3) (4) (5)
Share Free-State Peers (sd) 0.053∗∗ 0.054∗∗ 0.055∗∗∗ 0.056∗∗ 0.054∗∗
(0.021) (0.021) (0.021) (0.022) (0.022)
Slave Share (County) $\times$ Share Free-State Peers (sd) -0.036 -0.034 -0.036 -0.047∗∗ -0.041∗∗
(0.018) (0.018) (0.020) (0.021) (0.020)
Slave Share (County) -0.166∗∗∗ -0.164∗∗∗ -0.159∗∗∗ -0.142∗∗∗ -0.161∗∗∗
(0.033) (0.033) (0.034) (0.034) (0.034)
Baptists $\times$ Share Free-State Peers (sd) 0.005
Baptists -0.040
FriendQuaker $\times$ Share Free-State Peers (sd) 0.004
FriendQuaker 0.022
Modern Republican $\times$ Share Free-State Peers (sd) -0.013
Modern Republican 0.091∗∗
Southern Democrat $\times$ Share Free-State Peers (sd) -0.003
Southern Democrat -0.021
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 0.358 0.358 0.358 0.330 0.330
Observations 352 352 352 327 327
R-squared 0.341 0.347 0.342 0.349 0.337
Note. Our sample for this table is restricted to cadets from Slave States. All regressions include standardized versions of Slave share (county), Baptists, FriendsQuaker, Modern Republican, and Southern Democrat. We exclude 36 observations with missing county-level information. The control variables consist of Age in 1860, Class Rank, and Cohort. Columns (4)–(5) have smaller sample sizes because Washington, DC did not have complete voting records prior to 1861. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ B.5 Peer Influence by Cohort
B.6 Peer influence and continuous military service experience
Table S10: Peer composition and allegiance choice: continuous military service experience
@@ B.6 Peer influence and continuous military service experience
Non-continuous Military Service Continuous Military Service
@@
Share Free-State Peers (sd) 0.014 0.022 0.005 0.098∗∗∗ 0.063∗∗ 0.062∗∗
@@
Age in 1860 -0.006 0.013 -0.017 -0.014
(0.022) (0.024) (0.017) (0.018)
Class Rank -0.001 -0.001 0.002 0.002
(0.001) (0.001) (0.001) (0.001)
Slave Pop. Share (sd) -0.171∗∗∗ -0.218∗∗∗
(0.030) (0.024)
Cohort -0.011 0.006 -0.023 -0.020
(0.021) (0.022) (0.016) (0.017)
@@
Dependent var. mean 0.202 0.202 0.202 0.431 0.431 0.431
Observations 119 119 119 269 269 269
R-squared 0.001 0.230 0.347 0.039 0.272 0.323
Note. The sample for this analysis consists of all war-participating cadets. Columns (1)-(3) use the sample of cadets who remained out of military service for more than one year before the war, while Columns (4)-(6) use the sample of cadets who remained continuously in military service. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations ***$p<0.01$, **$p<0.05$, *$p<0.1$.